### v3

# Estimating Model Performance Under Covariate Shift Without Labels

Abstract
Machine learning models often experience performance degradation post-deployment due to shifts in data distribution. It is challenging to assess model’s performance accurately when labels are missing or delayed. Existing proxy methods, such as drift detection, fail to measure the effects of these shifts adequately. To address this, we introduce a new method, Probabilistic Adaptive Performance Estimation (PAPE), for evaluating classification models on unlabeled data that accurately quantifies the impact of covariate shift on model performance. It is model and data-type agnostic and works for various performance metrics. Crucially, PAPE operates independently of the original model, relying only on its predictions and probability estimates, and does not need any assumptions about the nature of the covariate shift, learning directly from data instead. We tested PAPE on tabular data using over 900 dataset-model combinations created from US census data, assessing its performance against multiple benchmarks. Overall, PAPE provided more accurate performance estimates than other evaluated methodologies.
1 Introduction
The final step in the machine learning (ML) model development process is to evaluate the model on an unseen dataset known as a test set. Successful performance on this dataset, indicating potential business value, typically leads to the model’s deployment to production. The expectation is that the model’s performance on the test set will be replicated with production data [1]. Yet, this assumption often fails due to data distribution shifts that lead to model performance degradation [2]. Therefore, ongoing monitoring of the model’s performance is essential to ensure it continues to meet expected outcomes. However, real-time performance assessment is challenging in many scenarios because true labels are either unavailable or significantly delayed.
In scenarios where labels are unavailable, performance stability is commonly estimated using proxy methods that primarily monitor changes in the input data distribution [3]. However, such changes in input data often have minimal effect on the actual performance of the model [4]. Additionally, the methods used to measure changes in the input distribution typically provide estimates in units that do not correspond to performance metrics. This results in at best a qualitative understanding of performance stability. When a distribution shift does occur, these methods do not provide any insight into the direction or magnitude of the impact on performance. The true extent of performance changes only becomes apparent once the labels become available, which may sometimes never happen. Until then, the model might be incurring significant financial losses [5].
Accurate performance estimates without labels can address this risk. With these estimates, the impact of changes in model performance on business operations can be quantified, allowing for informed decisions. These may include altering how model predictions are utilized in downstream processes, discontinuing their use altogether, or determining whether model retraining is required.
The main contributions of this paper are the following:
  • We introduce a novel algorithm called PAPE (Probabilistic Adaptive Performance Estimation) for assessing model performance without labels under covariate shift. PAPE enables estimation of various classification performance metrics and is designed to be both model and data-type agnostic. Importantly, it autonomously learns from data without requiring user input regarding the nature of the covariate shift.
  • We present theoretical and empirical analyses demonstrating how PAPE can effectively yield accurate performance estimations.
  • We develop an evaluation methodology specifically tailored to assess performance estimation methods. This includes a selection of relevant datasets, their preprocessing steps, an implementation of benchmarks suitable for tabular data and performance metrics of interest, and new evaluation metrics relevant to the model monitoring context.
  • Our extensive evaluations show that PAPE significantly outperforms existing benchmarks across datasets and metrics used.
In Section 2, we present the related work. In Section 3, we introduce the PAPE algorithm, presenting its implementation, theoretical foundations, guarantees, and limitations. Section 4 describes the evaluation methodology we used to assess performance estimation methods and presents the results. Section 5 summarizes the results and outlines future research directions.
2 Related work
Importance Weighting (IW) [6] is a simple but powerful method of estimating model performance. IW calculates a weighted performance metric by weighing the labeled reference data based on the likelihood ratio of observing them in the production distribution compared to the reference distribution. An extended version of IW [7] takes into account the nature of covariate shift and its potential impact on performance, as defined by a user. Confidence-based methods form another family of approaches. Confidence-based performance estimation (CBPE) [8] calibrates probability estimates using labeled reference data and calculates expected confusion matrix elements on production data to turn them into the desired performance metric. Average threshold confidence (ATC) [9] learns a threshold on confidence estimates and predicts accuracy using the fraction of predictions that exceed the threshold. Difference of confidence (DoC) [10] uses the difference between confidence estimates from test to production data and fits a regression model that maps this difference into the performance metric. DoC is an example of a distance-based approach that measures the distance between test and production data and identifies the relationship between the distance and the expected performance. Another example of this approach is using Fréchet distance and linear regression [11]. Other approaches, such as Contrastive Automated Model Evaluation (CAME) [12], use contrastive learning. Finally, reverse testing (RT) [13] methods use the monitored model predictions as labels to train a reverse model on production data. Then, they assume that the reverse model’s performance on the labeled reference dataset corresponds to the monitored model’s performance on the production data.
3 Probabilistic Adaptive Performance Estimation (PAPE)
In this section, we present PAPE algorithm. In  3.1, we describe the algorithm’s inputs, outputs, and steps. Section 3.2 gives a verbose description of the notation and provides proof for using the main formula in Algorithm 1.
3.1 The algorithm
The steps, required inputs, and outputs of the PAPE algorithm are detailed in Algorithm 1.
Algorithm 1 Probabilistic Adaptive Performance Estimation (PAPE)
Inputs
  • Labeled reference data: inputs $x_{ref}$, labels $y_{ref}$ and number of observations $n_{ref}$
  • Unlabeled production data: inputs $x_{prod}$ ; $n_{prod}$ observations
  • Uncalibrated probability estimates (scores) of binary classifier $f$ on reference and production data: $f(x_{ref})$, $f(x_{prod})$
  • Actionable predictions of $f$ on production data: $g\bigl{(}f(x_{prod})\bigr{)}$
  • Function $m$ that computes the performance metric: $m\bigl{(}g(f(x)),y\bigr{)}$
Outputs
  • Estimation of performance metric $m$ of classifier $f$ on unlabeled production data - $\widehat{m}_{prod}$
  1. 1.
    Estimate density ratios $w(x)=\frac{p(x)_{prod}}{p(x)_{ref}}$ for reference inputs $x_{ref}$:
    1. a.
      Assign label $z_{ref}=0$ to reference inputs $x_{ref}$ and $z_{prod}=1$ to production inputs $x_{prod}$
    2. b.
      Concatenate inputs $x_{rp}=\mathrm{conc}(x_{ref},x_{prod})$ and labels $z_{rp}=\mathrm{conc}(z_{ref},z_{prod})$
    3. c.
      Fit density-ratio-estimation classifier (DRE) on $(x_{rp},z_{rp})$
    4. d.
      Calculate probability estimates for reference data $\widehat{p}\left(z=1|x_{ref}\right)=\mathrm{DRE}\left(x_{ref}\right)$
    5. e.
      Estimate density ratios for reference data $\widehat{w}\left(x_{ref}\right)=\frac{n_{ref}}{n_{prod}}\frac{\widehat{p}(z=1|x_{ref})}{\widehat{p}(z=0|x_{ref})}$
  2. 2.
    Fit weighted calibrator (regressor) $c$ on reference data with inputs $f(x_{ref})$, targets $y_{ref}$, and weights $\widehat{w}(x_{ref})$
  3. 3.
    Use $c$ to get calibrated probabilities on production data: $c\bigl{(}f(x_{prod})\bigr{)}$
  4. 4.
    Estimate performance metric $m$ on production data with:
    $\widehat{m}_{prod}=\mathrm{mean}_{x\sim prod}\left[c\bigl{(}f(x)\bigr{)}m\bigl{(}g(f(x)),1\bigr{)}+\bigl{(}1-c(f(x))\bigr{)}m\bigl{(}g(f(x)),0\bigr{)}\right]$ (1)
3.2 Theoretical guarantees
Consider a binary classification task with inputs $x\in X$ and targets $y\in Y$, where $Y=\{0,1\}$. Fix a distribution $D$ and assume that $(x,y)\sim D$. Assume an uncalibrated classifier $f$ that for each observation $x\in X$ calculates probability estimates $f(x)=\hat{P}[y=1|x]$, referred to as scores $s$, so that $f:X\rightarrow S$, where $S=[0,1]$. Suppose that there is a function $c$ that calibrates scores $s$ into probabilities, that is $c(s)=P[y=1|f(x)=s]$ for all $s\in S$. Assume a function $g$ that maps scores returned by the classifier into actionable predictions. Consider a performance metric of interest $m$ that is calculated based on predictions of the classifier $g(f(x))$ and corresponding true labels $y$ that is $m\bigl{(}g(f(x)),y\bigr{)}$.
Lemma 1.
Performance metric $m$ of uncalibrated classifier $f$ on distribution $D$ can be estimated without using labels $y$ with:
$\widehat{m}_{(f,g,D)}=\mathrm{E}_{D}\left[c\bigl{(}f(x)\bigr{)}m\bigl{(}g(f(x)),1\bigr{)}+\bigl{(}1-c(f(x))\bigr{)}m\bigl{(}g(f(x)),0\bigr{)}\right]$ (2)
Proof.
Since all the expected values and probabilities are calculated over distribution $D$ we will omit the subscript $D$ for readability. We will additionally consider that $c$ calibrates scores $s$ only approximately, with expected absolute calibration error $\varepsilon$, that is:
$\mathrm{E}\bigl{[}y|f(x)=s\bigr{]}=P\bigl{[}y=1|f(x)=s\bigr{]}=c(s)+\epsilon_{s}$ (3)
$\sum_{s\in S}P\bigl{[}f(x)=s\bigr{]}E\Bigl{[}\bigl{|}y-c(s)|f(x)=s\bigr{|}\Bigr{]}=\sum_{s\in S}P\bigl{[}f(x)=s\bigr{]}\left|\epsilon_{s}\right|=\varepsilon$ (4)
We can now write:
$m_{\left(f,g\right)}$ $=\mathrm{E}\left[m\bigl{(}g(f(x)),y\bigr{)}\right]$
$=\sum_{s\in S}P\bigl{[}f(x)=s\bigr{]}\mathrm{E}\Bigl{[}m\bigl{(}g(f(x)),y\bigr{)}|f(x)=s\Bigr{]}$
$=\sum_{s\in S}P\bigl{[}f(x)=s\bigr{]}\mathrm{E}\Bigl{[}m\bigl{(}g(s),1\bigr{)}P\bigl{[}y=1|f(x)=s\bigr{]}+m\bigl{(}g(s),0\bigr{)}P\bigl{[}y=0|f(x)=s\bigl{]}\Bigr{]}$
$=\sum_{s\in S}P\bigl{[}f(x)=s\bigr{]}\mathrm{E}\Bigl{[}m\bigl{(}g(s),1\bigr{)}\bigl{(}c(s)+\epsilon_{s}\bigr{)}+m\bigl{(}g(s),0\bigr{)}\Bigl{(}1-\bigl{(}c(s)+\epsilon_{s}\bigr{)}\Bigr{)}\Bigr{]}$
$=\sum_{s\in S}P\bigl{[}f(x)=s\bigr{]}\mathrm{E}\Bigl{[}m\bigl{(}g(s),1\bigr{)}c(s)+m\bigl{(}g(s),0\bigr{)}\bigl{(}1-c(s)\bigr{)}\Bigr{]}$
$+\sum_{s\in S}P\bigl{[}f(x)=s\bigr{]}\mathrm{E}\Bigl{[}\epsilon_{s}\Bigl{(}m\bigl{(}g(s),1\bigr{)}-m\bigl{(}g(s),0\bigr{)}\Bigr{)}\Bigr{]}$
$=\widehat{m}_{(f,g)}+\sum_{s\in S}P\bigl{[}f(x)=s\bigr{]}\mathrm{E}\Bigl{[}\epsilon_{s}\Bigl{(}m\bigl{(}g(s),1\bigr{)}-m\bigl{(}g(s),0\bigr{)}\Bigr{)}\Bigr{]}$
If calibration is perfect, that is $\epsilon_{s}=0,\forall s\in S$, then the estimated metric equals the true metric. ∎
In other cases, performance estimation is as accurate as the calibration process. For example, for performance metrics for which $\max[m(g(s),1)-m(g(s),0)]=1$ the absolute error of the performance metric estimation is limited by the expected absolute calibration error:
$|m_{\left(f,g\right)}\ -\widehat{m}_{(f,g)}|\leq\sum_{s\in S}P[f(x)=s]|\epsilon_{s}|=\varepsilon$ (5)
To use equation (2) for estimating the performance of $f$ on unlabeled data from production distribution $D_{prod}$ we need $c$ a regressor that calibrates the scores on $D_{prod}$. We only have labels for the reference distribution to train the calibrator $c$ However, when probabilities are calibrated under reference distribution $D_{ref}$ and with respect to density ratios $w(x)_{ref\rightarrow prod}=\frac{p(x)_{prod}}{p(x)_{ref}}$ they become calibrated under production distribution [14]. That is true under the assumption that the true conditional probabilities of labels given the inputs remain the same for reference and production distribution, that is $p(y|X)_{ref}=p(y|X)_{prod}$. The input distribution may still change so that $p(x)_{ref}\neq p(x)_{prod}$. The probability density rates $w(x)_{ref\rightarrow prod}$ describe the covariate distribution shift from reference to production. Calibrator $c$ provides calibrated probabilities under $D_{prod}$ when:
$\mathrm{E}_{D_{ref}}\bigl{[}w(x)_{ref\rightarrow prod}\cdot y|f(x)=s\bigr{]}=c\bigl{(}f(x)\bigr{)},\forall s\in S$ (6)
3.3 Limitations
The main assumption of PAPE is that the probability distribution of labels conditioned on inputs remains the same between reference and production data, that is $p(y|X)_{ref}=p(y|X)_{prod}$. Even if that holds, the method will not work when data shifts to a region that is not covered by the reference distribution. That is when there is a region $X_{r}$ for which $p(X_{r})_{prod}>0$ but $p(X_{r})_{ref}=0$. Weighted calibration will not be possible for observations coming from that region. Another issue is that PAPE uses equation 1 that assumes that the calibrator $c$ provides calibrated probabilities, but that is rarely the case when working on finite samples of data (due to both - difficulties in learning accurate $w$ weights and the calibration process itself). Even with perfectly calibrated probabilities, the realized sample performance may end up different from the estimated one due to uncertainty related to probabilities realizing into binary targets. Even though PAPE algorithm is data-type agnostic, the DRE estimator will require different implementation for data types other than tabular data. Finally, equation 1 can be used directly to estimate a performance metric that can be defined as a mean of observation-level performance values, like accuracy. For confusion-matrix-based metrics, like F1, confusion matrix elements needs to be estimated first and then the metric of interest can be approximated. Lastly, standard concerns about using machine learning algorithms apply, such as the consequences of learning from finite samples of data.
4 Evaluation Methodology
This section describes our evaluation methodology. We describe the datasets we used in 4.1, how we created the ML models whose performance was estimated, and the experimental setup used for performance estimation evaluation in 4.2. We provide a description of the evaluated methods in 4.3. Section 4.4 shows and discusses the results. Additionally, we analyze sample size effects in 4.5.
4.1 Datasets
The datasets we use to evaluate the methods were derived with Folktables [15]. Folktables uses US Census data and preprocesses them to create a set of binary classification problems. We use the following tasks: ACSIncome, ACSPublicCoverage, ACSMobility, ACSEmployment, ACSTravelTime. For each of the five prediction tasks listed above, we fetch Folktables data for all 50 US states spanning four consecutive years (2015-2018). This gives us 250 datasets.
4.2 Experimental setup
First, we fit machine learning models whose performance will be estimated. We separate the first-year data (2015) in each fetched dataset and use it as a training period. We fit 5 commonly used binary classification algorithms on the training data: Logistic Regression [16], neural network model [17], Random Forest [16], XGBoost [18], and LGBM [19] with default parameters. We use these models to predict labels on the remaining part of the dataset. We record both - binary predictions and probability estimates. We end up with 1250 dataset-model pairs, which we call evaluation cases.
The rest of the data for each case is further divided into two periods - reference (the year 2016) and production (2017, 2018). Reference data is where the performance estimation methods are fitted using the monitored model inputs, outputs, and targets. Production data is further split into data chunks of 2000 observations each, maintaining the order of the observations. The realized performance of the monitored model is calculated on each chunk based on labels and monitored model outputs. The performance of the monitored model on each data chunk is also estimated based on monitored model inputs and outputs. The realized performance and the estimated performance are compared - each production chunk is an evaluation point of the performance estimation methods. After applying filtering that removed evaluation cases with less than 6000 observations (3 chunks) in the reference set and cases where the performance of the monitored model on the reference data was worse than random (with AUROC lower than 0.5 or F1 equal 0), we ended up with 959 evaluation cases. These contain 36 675 production data chunks (evaluation points) for each evaluated method. The metrics estimated are accuracy, AUROC, and F1. Figure 1 shows an example of an AUROC estimation result.
Figure 1: Estimation of AUROC for ACSIncome task data from California and LGBM as the monitored model. The black line is the realized AUROC of the monitored model. The red line and brown dashed line are estimates respectively from PAPE and the TEST SET performance benchmark.
We evaluate the performance estimation methods listed in 4.3 on all the production chunks in all the appropriate evaluation cases [^1] We ran the experiments on a single Intel i7-11800H 2.30GHz, 32GB RAM machine. Data preparation took about 48 hours (partially limited by internet bandwidth) while evaluating the methods took over 29 hours.. The evaluation metrics are described, and the results are shown in 4.4.
4.3 Evaluated methods
Below, we briefly describe the evaluated methods. We omit MANDOLINE from [7] as it requires user input on the nature of covariate shift.
PAPE
We implement PAPE following algorithm 1. In step 1 we estimate the weights $w$ with DRE as described in [20, chapter 2.7.5]. We use LGBM Classifier with default hyperparameters for this. In steps 2 and 3, we approximate weighted calibration under reference distribution and weights $w$ by fitting $w$-weighted regressor on reference data. We use LGBM Regressor with default hyperparameters.
TEST SET performance
For each evaluation data chunk from production data, the estimated performance equals the performance calculated on reference data (and typically, the test set is chosen as the reference set, hence the name). It represents the initial assumption under which the model’s performance on the production data is constant and equal to the performance calculated on the test set.
Average Threshold Confidence (ATC)
With ATC, we take the raw scores provided by the monitored model $f(x)$ and apply the maximum confidence function to them, that is: $\mathrm{MC\bigl{(}f(x)\bigr{)}}=f(x)$ when $f(x)\geq 0.5$ and $1-f(x)$ otherwise. Then, we use it to learn a threshold on reference data such that the fraction of observations above the threshold equals the performance metric value calculated on reference data. When inferring, we apply MC to the evaluation chunk and calculate the fraction of observations above the learned threshold. The estimated metric is equal to the calculated fraction.
Difference of Confidence (DoC)
DoC was proposed to estimate image classifier accuracy [10]. The method assumes that the performance change is proportional to the change in the mean of maximum confidence. To learn this relationship, we fit Linear Regression on the mean difference of confidence between the shifted and the reference data as inputs and the difference in performance between these two as the targets. In order to get datasets with synthetic distribution shifts we perform multiple random re-samplings of the reference dataset.
Modified Reverse Testing (RT-mod)
Reverse Testing methods train a reverse model on production data with the monitored model inputs as features and monitored model predictions as targets. This reverse model is then used to make predictions on reference inputs, and its performance is evaluated with the reference labels. The reverse model’s performance on reference data is a proxy for the monitored model’s performance on the production data. We modified the method by finding a constant to ensure that the average performance estimate from the reverse model on reference data matches the realized performance. Without this modification the estimates returned by the method were always biased - either over or underestimating the performance metric.
Importance Weighting (IW)
For IW, we first estimate density ratios $\widehat{w}(x)_{ref\rightarrow prod}$ between reference data and the production data chunks with the DRE classifier - the same we use for PAPE. Then, we use them as weights to calculate the weighted performance metric of interest using reference data.
Confidence-based Performance Estimation (CBPE)
CBPE is a simpler version of PAPE. The difference is that CBPE calibrates on reference data without reweighing to production distribution, so it does not fully account for covariate shift impact.
4.4 Results
Performance estimation is a regression problem. This leads to evaluation metrics such as mean absolute error (MAE) or root mean squared error (RMSE). However, the MAE/RMSE of all data chunks from multiple evaluation cases is meaningless. Large MAE/RMSE for an evaluation case where the performance of the model has a large variance might still be satisfactory. On the other hand, even small changes (in the absolute scale) in performance might be significant for cases where the model performance is very stable. To account for this, we scale absolute/squared errors by the standard error (SE) calculated for each evaluation case and performance metric. We first measure SE as the standard deviation of the realized performance sampling distribution on the reference data. We do this by randomly sampling with repetitions 2000 observations (the size of the evaluation chunk) from the reference data and calculating the realized performance metric on each sample. We repeat this 500 times and calculate the standard deviation of the obtained per-sample metrics - the standard error. Then, we divide absolute and squared errors by SE and squared SE, respectively. We call the SE-scaled metrics mean absolute standard error (MASTE) and root mean squared standard error (RMSSTE) (7).
$\mathrm{MASTE}$ $=\frac{1}{\sum^{k}_{i=1}c_{i}}\sum\limits_{i=1}^{k}\sum\limits_{j=1}^{c_{i}}\frac{\left|m_{i,j}-\widehat{m}_{i,j}\right|}{\mathrm{SE}_{i}}$ (7)
$\mathrm{RMSSTE}$ $=\sqrt{\frac{1}{\sum^{k}_{i=1}c_{i}}\sum\limits_{i=1}^{k}\sum\limits_{j=1}^{c_{i}}\frac{\left(m_{i,j}-\widehat{m}_{i,j}\right)^{2}}{\mathrm{SE}_{i}^{2}}}$
where $i$ is the evaluation case index ranging from $1$ to $k$, and $j$ is the index of the production data chunk ranging from $1$ to $c_{i}$ (number of production chunks in a case $i$). $m_{i,j}$ and $\widehat{m}_{i,j}$ are respectively realized and estimated performance for case i and chunk j. $\mathrm{SE}_{i}$ is the standard error of the $i$-th evaluation case. The results are shown in Table 1.
Table 1: MASTE and RMSSTE of the evaluated methods for each estimated metric.
Accuracy AUROC F1 MASTE RMSSTE MASTE RMSSTE MASTE RMSSTE TEST SET 1.62 2.88 1.45 2.30 2.53 8.23 RT-mod 2.31 5.41 1.85 4.29 2.06 4.74 ATC 1.58 2.79 1.90 3.52 2.97 9.06 DOC 1.13 1.80 1.37 2.75 2.56 8.52 CBPE 1.08 1.75 1.07 1.68 1.03 2.12 IW 1.04 1.40 1.06 1.56 1.07 1.74 PAPE 0.97 1.28 0.99 1.45 0.90 1.34
PAPE, together with CBPE and IW, shows strong improvement over the TEST SET baseline. PAPE significantly outperforms the second-best method for each estimation and evaluation metric (paired t-test [21] p-value $=0.00$). PAPE result of MASTE equal to 0.97 for accuracy estimation means that the estimation is, on average, less than 1 SE away from the realized performance. For intuition, the MASTE of PAPE for the evaluation case shown in Figure 1 is equal 1.11. The TEST SET performance MASTE for accuracy is 1.62, indicating that accuracy changes significantly in the production chunks. If the performance was stable with only random normal fluctuations, the MASTE of the TEST SET performance would be around 0.8. That is because the average absolute deviation for normal distribution is $\sqrt{\frac{2}{\pi}}\approx 0.8$ of its standard deviation. The changes for F1 are much stronger, as shown by the increased MASTE of TEST SET - 2.53. This explains why performance estimation methods provide the strongest improvement compared to the TEST SET performance for this metric.
We also analyze the quality of estimations across different levels of realized performance changes. The TEST SET performance baseline is accurate for relatively small realized performance changes (within +/- 2SE). Other methods should match or exceed this baseline, particularly when realized performance changes are significant. We sorted the data by absolute performance change magnitude and calculated rolling MASTE for 2SE-wide intervals. The results are shown in Figure 2.
Figure 2: Quality of performance estimation [MASTE of estimated metric] vs. absolute performance change [SE]. The x axis indicates the center of the data bucket - for example, value 1 indicates a bucket that contains data chunks for which the absolute performance change was in between 0 - 2 SE. The left y axis shows MASTE of the evaluated method for the data bucket. The right y axis shows the number of data chunks in each bucket on a logarithmic scale. Methods not included in the plots (like RT-mod in plot A) performed significantly worse. Including them would change the scale of the plot and make comparison between best-performing solutions difficult.
For all the estimated metrics, PAPE estimation error is the lowest in nearly whole evaluated region. IW and CBPE show comparable accuracy.
4.5 Sample size effect
The experiments described above were run for arbitrarily chosen size of the evaluation data chunk (2000 observations), so we have also investigated the quality of performance estimates for different chunk sizes. We split production data into chunks of the following sizes: 100, 200, 500, 1000, 2000, and 5000. The step, that is the difference between the beginning of each chunk, was set to 1000 observations for each chunk size. Due to computational complexity, we ran the experiment only for one evaluation case  [^2] Computation took over 9 hours.. We selected the biggest data set (California), the prediction task for which performance changes significantly (ACSEmployment), the default choice algorithm for the classification task on tabular data - LGBM, and the default metric - AUROC. Since the results come from a single evaluation case, we used mean absolute error (MAE). The results in Figure (3) show that all methods give less accurate AUROC estimates for smaller chunk sizes. This is expected as the random noise effects are more significant for small samples. For the evaluated case, this has the strongest impact on IW accuracy of estimation.
Figure 3: Effect of sample size on mean absolute error of AUROC estimation. Calculated for sample sizes of 100, 200, 500, 1000, 2000, and 5000, on data from California, prediction task ACSEmployment, client model: LGBM.
5 Conclusions
We introduced PAPE, a novel approach for accurately estimating the performance of classification models without labels when faced with covariate shift. We developed an evaluation methodology for performance estimation methods. This methodology encompasses dataset preparation, benchmark implementation, and the introduction of new evaluation metrics that are essential for aggregating results from multiple datasets and model monitoring scenarios. We employed this evaluation methodology to assess PAPE alongside other methods, analyzing over 900 model-dataset pairs and generating more than 36,000 data chunks for thorough evaluation. The results show that PAPE provided more accurate performance estimates than other evaluated methods.
Looking ahead, our future work will focus on quantifying the errors and uncertainties associated with these estimates and developing estimation intervals. A key source of error in our current methodology is the neglect of calibration errors, as indicated by PAPE’s use of equation 1. To address this, we plan to implement the calibration of $f$ within a cross-validation loop and estimate the conditional target probability $P(Y=1|s)$ for calibrated scores as a probability distribution, specifically using a Beta distribution. This will allow us to sample from this distribution for each $s$ and generate, using equation 1, a distribution of estimated performance metrics. Apart from that, we are actively looking for other sources of relevant datasets to extend our evaluation and better assess performance estimation methods.
Acknowledgments and Disclosure of Funding
The research described in this paper was done within NannyML - a venture-backed open-source software company focusing on post-deployment data science solutions. The company has received public R&D funding from Flanders Innovation & Entrepreneurship (VLAIO) - project number: HBC.2022.0846.
{CJK*}
UTF8gbsn
Appendix A Appendix / Estimating metrics other than accuracy
Let us recall equation (1) of the PAPE algorithm that allows to estimate performance metric $m$ on production data based on monitored model predictions $g\bigl{(}f(x)\bigr{)}$ and calibrated probabilities $c\bigl{(}f(x)\bigr{)}$:
$\widehat{m}_{prod}=\mathrm{mean}_{x\sim prod}\left[c\bigl{(}f(x)\bigr{)}m\bigl{(}g(f(x)),1\bigr{)}+\bigl{(}1-c(f(x))\bigr{)}m\bigl{(}g(f(x)),0\bigr{)}\right]$
This equation can be used directly to estimate a performance metric that can be calculated as a mean of observation-level values, like accuracy. For metrics based on confusion matrix elements like F1, precision, recall and other, we first estimate these elements’ expected values. Generally function $g$ is any transformation that turns scores $f(x)$ into actionable predictions. For performance metrics based on confusion matrix $g$ needs to return binary predictions - one of which is to be considered positive and the other - negative. Typically $g\bigl{(}f(x)\bigr{)}$ returns 00 or $1$ and we consider these predictions, respectively, negative and positive. Assuming that is the case we estimate confusion matrix elements with:
$\widehat{tp}_{prod}$ $=\mathrm{mean}_{x\sim prod}\Bigl{(}c\bigl{(}f(x)\bigr{)}\cdot g\bigl{(}f(x)\bigr{)}\Bigr{)}$
$\widehat{fp}_{prod}$ $=\mathrm{mean}_{x\sim prod}\Bigl{(}\Bigl{(}1-c\bigl{(}f(x)\bigr{)}\Bigr{)}\cdot g\bigl{(}f(x)\bigr{)}\Bigr{)}$
$\widehat{tn}_{prod}$ $=\mathrm{mean}_{x\sim prod}\Bigl{(}c\bigl{(}f(x)\bigr{)}\cdot\Bigl{(}1-g\bigl{(}f(x)\bigr{)}\Bigr{)}\Bigr{)}$
$\widehat{fn}_{prod}$ $=\mathrm{mean}_{x\sim prod}\Bigl{(}\Bigl{(}1-c\bigl{(}f(x)\bigr{)}\Bigr{)}\cdot\Bigl{(}1-g\bigl{(}f(x)\bigr{)}\Bigr{)}\Bigr{)}$
Having these we can, for example, estimate recall with:
$\widehat{\mathrm{recall}}_{prod}\approx\frac{\widehat{tp}_{prod}}{\widehat{tp}_{prod}+\widehat{fn}_{prod}}$
To calculate (and estimate) AUROC we do not need model binary predictions $g\left(f(x)\right)$ instead we use raw scores $f(x)$ returned by the classifier. We define a new function $h_{t}$:
$h_{t}\bigl{(}f(x),t\bigr{)}=\begin{cases}1,&f(x)\geq t\\ 0,&f(x)<t\end{cases}$
Using this function we estimate confusion matrix elements for all distinct scores returned by the classifer, that is for all $t\in S$. Having those we estimate true positive rates (TPR) and false positive rates (FPR). Finally, we calculate the area under the curve created from estimated TPRs and FPRs to get AUROC.
Notice that the formula for recall gives only approximate result. The same applies to TPRs and FPRs obtained in the described way. That is because in general case the quotient of expected values of the random variables is not equal to the expected value of the quotient. In order to get an actual expectation of confusion-matrix-based metric we could perform sampling, create its distribution and calculate the expected value. We did not do this because it is computationally expensive, and the approximate formula gave accurate enough results for the metrics we experimented with (F1 and AUROC). In future research we aim to address this problem by providing efficient solution to calculate unbiased expectation of the performance metrics of interest.